OpenStudy (hiimky):

Write an equation to model the distance between the point (2,4) and any point along the curve y=4x^3+1

OpenStudy (lord_box):

You can accomplish this using the distance formula.

OpenStudy (lord_box):

\[d = \sqrt(x_2 - x_1)^2 + (y_2 - y_1)^2\]

OpenStudy (hiimky):

How would I put the equation into the distance formula?

OpenStudy (hiimky):

could i solve in terms of each variable

OpenStudy (lord_box):

Well when y = f(x), you can plug in f(x) in for one of the y.

OpenStudy (lord_box):

\[d = \sqrt(x - 2)^2 + (f(x) - 4)^2\]

OpenStudy (hiimky):

Thank you for your help